What is the gram solubility of CAF2 in 100. mL of pure water at 25°c?
The solubility product constant of CAF2 at 25°c is 3.9x10^-11
Answer #1

caf, a ca2t + 2F- kep = [Ca 2+] [f-] ² Let sbe malar Salubility kep = (S) (28)2 = 483 kep= molar Solubility 3.9x10-1 = 483 co